two cars that are traveling toward each other are 455 miles apart. One car is traveling 10 miles per hour faster than the other car. The cars pass after 3.5 hours. How fast is each car traveling?
Answer by checkley77(12844)

D=RT
455=(R+R+10)3.5
455=(2R+10)3.5
455=7R+35
7R=455-35
7R=420
R=420/7
R=60 MPH IS THE RATE OF THE SLOWER CAR.
60+10=70 MPH IS THE RATE FOR THE FASTER CAR.
PROOF:
455=(60+70)*3.5
455=130*3.5
455=455